OpenStudy (anonymous):

Find the coordinates of the center and the length of the radius of the following circle. (x - 3)^2 + (y - 7)^2 = 49

OpenStudy (anonymous):

(3,7) , 7

OpenStudy (anonymous):

(3,-7) r=7

OpenStudy (anonymous):

(3,7) radius 7

OpenStudy (ash2326):

standard equation of a circle (x-a)^2+(y-b)^2=r^2 center (a, b) radius r compare it with the given equation center (3,7) and radius 7

OpenStudy (anonymous):

(x-a)^2 + (y-b)^2 = r^2 centre is (a,b), radius = r